# Explain What You Mean: Intent Augmented Knowledge Graph Recommender Built With An LLM

### v2

@@ 3.3. Knowledge Graph Convolution Layer
Similar to KGCN (Wang et al., 2019), we employ a knowledge graph convolution layer to capture structural proximity among entities in a knowledge graph. The model learns node embeddings $\bm{E}\in R^{T_{R}\times d}$ and relation embeddings $\bm{R}\in R^{N\times d}$, where $T_{R}$ is number of relation types, $N$ is number of nodes, and $d$ is the embedding dimensionality. Denote the input features for some node $v$ as $\mathbf{v}$, the set of entities directly connected to $v$ as $\mathcal{S}_{v}$, then the output embedding is the summation-aggregated neighborhood entity and relation embeddings:
(3) $\mathbf{v}^{\text{out}}=\sigma\left({\bf W}\cdot[{\bf v}+\text{softmax}(\bm{R}[\mathcal{S}_{v}]\bm{E}[\mathcal{S}_{v}]^{T})\bm{E}[\mathcal{S}_{v}]]+{\bf b}\right)$
@@
3.4. Knowledge Graph Embedding Translation
@@
Due to the incomplete nature of KGs, KG completion is often leveraged as a self-supervised learning task, which predicts the missing entity $e_{h}$ or $e_{t}$ for a triplet $(e_{h},e_{t},r)$. To this end, TransE (Bordes et al., 2013), a popular knowledge graph embedding model commonly used for KG completion, enforces a translation in the embedding space: $\mathbf{e}_{h}+\mathbf{e}_{r}\approx\mathbf{e}_{t}$. This embedding translation is achieved via the following training objective: $\min\mathop{\Sigma}\limits_{\mathbf{e}_{h},\mathbf{r},\mathbf{e}_{t}}\parallel\mathbf{e}_{h}+\mathbf{r}-\mathbf{e}_{t}\parallel$.
@@
One issue with TransE embedding is that a single relation type may correspond to multiple head entities or tail entities, leading to significant 1-to-N, N-to-1, and N-to-N issues (Wang et al., 2014). As an improvement, TransH (Wang et al., 2014) learns different representations for an entity conditioned on different relations. It assumes that each relation owns a hyperplane, and the translation between head entity and tail entity is valid only if they are projected to the same hyperplane. It defines an energy score function for a triplet as follows:
(4) $f(\mathbf{e}_{h},\mathbf{e}_{t},\mathbf{r})=\parallel\mathbf{e}_{h}^{\bot}+\mathbf{r}-\mathbf{e}_{t}^{\bot}\parallel$
@@
(5) $\mathbf{e}_{h}^{\bot}=\mathbf{e}_{h}-\mathbf{w}_{r}^{\mathrm{T}}\mathbf{e}_{h}\mathbf{w}_{r}$
(6) $\mathbf{e}_{t}^{\bot}=\mathbf{e}_{t}-\mathbf{w}_{r}^{\mathrm{T}}\mathbf{e}_{t}\mathbf{w}_{r}$
@@
Finally, the training of TransH encourages the discrimination between valid triplets and incorrect ones using margin-based ranking loss. Similar to KTUP (Cao et al., 2019), we employ:
(7) $\mathcal{L}_{KG}=-\sum_{\begin{subarray}{c}(\mathbf{e}_{h},\mathbf{e}^{+}_{t},\mathbf{r}^{+})\in\mathcal{KG}\\ (\mathbf{e}_{h},\mathbf{e}^{-}_{t},\mathbf{r}^{-})\in\mathcal{KG}^{-}\end{subarray}}\log\sigma[f(\mathbf{e}_{h},\mathbf{e}^{-}_{t},\mathbf{r}^{-})-f(\mathbf{e}_{h},\mathbf{e}^{+}_{t},\mathbf{r}^{+})]$
@@
4. Methodology
In this section, we address the following key questions regarding LLM-based knowledge graph augmentation: Q1: How to enable node-level intent entity extraction with RAG? Q2: How to inject the learned intents as structural priors into KG modeling? Q3: How to improve scoring with intent awareness at test time?
4.1. Intents extraction with RAG (Q1)
In order to address challenges posed by sparse user behavior scenarios, we leverage the LLM to perform data augmentation. Unlike existing LLM recommendation methods that directly synthesize user-item interactions (Wei et al., 2024), we take a more conservative approach, tasking the LLM with a simpler, more reliable role to minimize noise in the generated outputs. This approach leverages the LLM’s pre-trained common sense knowledge for accurate distillation.
We note that a forward call to the LLM with appropriate instructions is able to extract an intent entity set. For example, for item node $i$, we denote the set of intent entities as $\Omega_{i}$, $\Omega_{i}=LLM($###Item Description: {Item i description} Return a list of entities mentioned in the Item Description that the user may have intents to interact with)))). However, there are two gaps of this naive application: (1) the existence of domain-specific knowledge may make it difficult for the LLM to fully understand the context; (2) the connectivity of the extracted intents could follow a very long-tail distribution, meaning that the majority of intent nodes might only link to few user/item nodes. To mitigate these challenges, we propose the following improvements over the naive forward call.
4.2. In-context Learning
To keep the resource budget lightweight, we curate an in-context learning process without fine-tuning. For the enterprise use-case, which differs from the public dataset use-case, user profiles and item descriptions often contain abbreviations and domain-specific concepts unknown to LLMs. To address this knowledge gap, we curate a knowledge base of key-value pairs, where keys are the domain-specific abbreviations/concepts, and values are their explanations. We identify any such terminology present in the user profile or item description and append their corresponding explanations to the prompt under a ###$\#\#\#Concept\ Explanation$ tag. This enriches the context for the LLM, enabling better understanding without the need for fine-tuning.
4.3. Intent connectivity enrichment via RAG
A two-round process extracts intents and builds a knowledge graph. In the first round, a simple prompt template is used to generate specific entities, linking users and items to intent nodes. However, this initial graph may be sparsely connected, with many intent nodes linked to only a few users/items. To address this, a second round enriches connectivity by linking additional user/item nodes to existing intent nodes, avoiding the computationally expensive alternative of grouping similar intents ($\mathcal{O}(N^{2})$ complexity)
In the second round, each user and item node is connected to additional intent entities from the fixed pool of intents generated in the first round. We call these new connections ”related intents”. During this second round of related intent selection, the existing intents for a given user/item node are excluded from the retrieval step.
Denote the intent entities obtained after the first round extraction as $\bar{\Omega}$. Given a user profile or an item description $text$, the second round intent extraction prompt can be formulated as: $LLM($###Knowledge Context: … ### Options: $R(x,\mathcal{N}\backslash\Omega_{x},K)$ What are the intents mentioned in $x$ that are the most relevant?)))) , where $R(x,\bar{\Omega}\backslash\Omega_{x}^{\prime},K)$ retrieves a group of $K$ intents that are semantically similar to the input text $x$, yet have not been extracted during the first round output ($\Omega_{x}^{\prime}$).
Figure 3. The graph module of IKGR. After constructing the knowledge graph, the embeddings of user, item and intents are fed into the knowledge graph convolution layer followed by an embedding translation layer.
4.4. Intent Translation GNN (Q2)

### v3

@@ 3.2 Knowledge Graph Convolution Layer
Similar to KGCN [27], we employ a knowledge graph convolution layer to capture structural proximity among entities in a knowledge graph. The model learns node embeddings $\bm{E}\in R^{N\times d}$ and relation embeddings $\bm{R}\in R^{T_{R}\times d}$, where $T_{R}$ is number of relation types, $N$ is number of nodes, and $d$ is the embedding dimensionality. Denote the input features for some node $v$ as $\mathbf{v}$, the set of entities directly connected to $v$ as $\mathcal{S}_{v}$, then the output embedding is the summation-aggregated neighborhood entity and relation embeddings:
$\mathbf{v}^{\text{out}}=\sigma\left({\bf W}\cdot[{\bf v}+\text{softmax}(\bm{R}[\mathcal{S}_{v}]\bm{E}[\mathcal{S}_{v}]^{T})\bm{E}[\mathcal{S}_{v}]]+{\bf b}\right)$ (3)
@@
4 Methodology
The key components of IKGR involve grounded user/item intent extraction with RAG, KG densification and a GNN prediction layer.
4.1 Intents Extraction with RAG
In order to address challenges posed by sparse user behavior scenarios, we leverage the LLM to perform data augmentation. Unlike existing LLM recommendation methods that directly synthesize user-item interactions [19], we take a more conservative approach, tasking the LLM with a simpler, more reliable role to minimize noise in the generated outputs. This approach leverages the LLM’s pre-trained common sense knowledge for accurate distillation.
We note that a forward call to the LLM with appropriate instructions is able to extract an intent entity set. For example, for item node $i$, we denote the set of intent entities as $\Omega_{i}$, $\Omega_{i}=LLM($###Item Description: {Item i description} Return a list of entities mentioned in the Item Description that the user may have intents to interact with)).
4.2 Factual Knowledge Access for Grounded Extraction
The intent extraction step assumes a high quality Item Description paragraph describing the intent. However, the Item Description can often be poorly-formulated in the datasets, leading to a gap in properly extracting intent. (i) In the case of enterprise search, the existence of private enterprise knowledge makes it difficult for the LLM to fully understand the context, and (ii) in the open sourced datasets case, only the item names are available without appropriate description - though details of these items are often available via online search or baked into the LLM, so that an agent with web access is able to augment it easily [28].
We bridge this gap by feeding the intent extraction module with enterprise private knowledge or open-world knowledge. Specifically, for our adapted enterprise use-case, user profiles and item descriptions often contain abbreviations and domain-specific concepts unknown to LLMs. We curate a knowledge base of key-value pairs, where keys are the domain-specific abbreviations/concepts, and values are their explanations. We identify any such terminology present in the user profile or item description and append their corresponding explanations to the prompt under a ###ConceptExplanation\#\#\#Concept\ Explanation tag. This enriches the context for the LLM, enabling better understanding without the need for fine-tuning. For open source datasets, the item name is first expanded into a paragraph summary by an LLM agent before it feeds into the entity extraction module.
4.3 Intent Connectivity Enrichment via RAG
The connectivity of the extracted intents could follow a very long-tail distribution, meaning that the majority of intent nodes might only link to few user/item nodes. To mitigate these challenges, a two-round process is used to extract and densify the intents KG. In the first round, a simple prompt template is used to generate specific entities, linking users and items to intent nodes. However, this initial graph may be sparsely connected, with many intent nodes linked to only a few users/items. To address this, a second round enriches connectivity by linking additional user/item nodes to existing intent nodes, avoiding the computationally expensive alternative of grouping similar intents ($\mathcal{O}(N^{2})$ complexity)
In the second round, each user and item node is connected to additional intent entities from the fixed pool of intents generated in the first round. We call these new connections "related intents". During this second round of related intent selection, the existing intents for a given user/item node are excluded from the retrieval step.
Denote the intent entities obtained after the first round extraction as $\bar{\Omega}$. Given a user profile or an item description texttext, the second round intent extraction prompt can be formulated as: LLM(LLM(###Knowledge Context: … ### Options: $R(x,\mathcal{N}\backslash\Omega_{x},K)$ What are the intents mentioned in $x$ that are the most relevant?)) , where $R(x,\bar{\Omega}\backslash\Omega_{x}^{\prime},K)$ retrieves a group of $K$ intents that are semantically similar to the input text $x$, yet have not been extracted during the first round output ($\Omega_{x}^{\prime}$). The full prompt is provided in Appendix A.
4.4 Generating Recommendation Candidates with Graph Module
After the KG has been built, a graph module is used to generate the recommendation candidate list. While multiple combinations of GNN architectures and loss functions could fit, we find a simple translation layer based architecture [29] that injects the learned intents as structural priors outperforms vanilla GNN options. We leverage this intent prior GNN as the default option in the experiment sections, and discuss the details in Appendix B. We also benchmark across three other options and show results in Section 5.3.
5 Experiments
We use Llama-3.1-8B model for LLM inference, and all-mpnet-base-v2 from sentence transformer for encoding textual features. We retrieve the top 100 prebuilt intent candidates in the kNN retrieval step of RAG. We apply an 8:1:1 ratio when sampling the positive/negative edges for train/dev/test sets in the graph.
[…]
Datasets Search Beauty Books Steam Yelp2022
#Users 36 03336\,033 40 22640\,226 251 394251\,394 281 428281\,428 1 987 8981\,987\,898
#Items 872 678872\,678 54 54254\,542 25 60625\,606 13 04413\,044 150 347150\,347
#Inter 3.5M 0.35M 3.2M 3.5M 6.9M
Density 0.011% 0.02% 0.05% 0.095% 0.002%
#Intents 495 285495\,285 39 30539\,305 45 93245\,932 53 94053\,940 65 04065\,040
#IntEdges 6.9M 209K 390K 231K 409K
AvgIntDeg 13.9 5.3 8.5 4.3 6.3
5.1 Baselines and Datasets
Enterprise Search. In the enterprise search setting, a set of developers query the search engine to search for datasets published by other developers within the enterprise. The search engine first uses BM25 [30] to retrieve a list of item candidates, then the IKGR is applied to rerank the search results. Evaluations are done over user’s feedback on how high the item of interest could be ranked among the final reranked list. In this setting, the items are enterprise datasets, which consist of text features such as dataset name, description, column names, ID labels, etc. The users are all registered within the enterprise and their developer profiles could be collected through separate channels to offer a hint about their dataset consumption preference. The IKGR is trained over historical user-dataset consumption interactions collected from separate channels. This dataset is labeled as Search in dataset description table and result tables.
Books, Beauty. These datasets are obtained from Amazon review [^1] http://jmcauley.ucsd.edu/data/amazon/ in [31], which contains a variety of categories. We utilize the Books and Beauty categories. We leverage the features of title, sales type, sales rank, categories, price, and brand.
Steam [^2] https://cseweb.ucsd.edu/~jmcauley/datasets.html#steam_data. This is a dataset collected from Steam [32], a large online video game distribution platform. We leverage the item features of app name, genres, publisher, sentiment, specs, tags.
To evaluate the ranking list of all the models, we employ Hit Ratio (HR), Normalized Discounted Cumulative Gain (NDCG), and Mean Reciprocal Rank (MRR). Considering we only have one ground truth item for each user, HR@$k$ is equivalent to Recall@$k$ and proportional to Precision@$k$; MRR is equivalent to Mean Average Precision (MAP). We report HR and NDCG with k=1,5,10k={1,5,10}. For all these metrics, the higher the value, the better the performance.
To verify the effectiveness of our method, we compare it with the following representative baselines. KGIN [13], which extracts non-readable intent in the latent embedding space. CSRec [10], which adds new nodes into KG with LLM, then post align the graph. HAKG [14], a hierarchy KG built upon the hyperbolic space. LLMRec [19], an LLM as interaction synthesizer approach. RippleNet [7], which uses a single layer of embedding translation by applying transH to user-item interaction. KGCN [27], KG based graph convolutional networks that captures inter-item relatedness effectively by mining their associated attributes on the KG. KTUP [29], which uses a single layer of embedding translation by applying TransH to user-item interaction.
Table 2: Performance comparison of different methods. Bold scores are the best in each row, while underlined scores are the second best.
Datasets Metric KGIN CSRec HAKG LLMRec RippleNet KGCN KTUP IKGR
Search HR@1 0.0074 0.0079 0.0080 0.0075 0.0024 0.0051 0.0078 0.0086
HR@5 0.0162 0.0156 0.0187 0.0158 0.0118 0.0161 0.0166 0.0202
HR@10 0.0263 0.0258 0.0255 0.0253 0.0218 0.0232 0.0262 0.0267
NDCG@5 0.0142 0.0144 0.0139 0.0148 0.0068 0.0101 0.0137 0.0151
NDCG@10 0.0154 0.0153 0.0161 0.0164 0.0087 0.0123 0.0142 0.0172
MRR 0.0135 0.0136 0.0128 0.0143 0.0091 0.0100 0.0130 0.0153
Beauty HR@1 0.1103 0.1401 0.1623 0.1563 0.0532 0.0984 0.1783 0.1369
HR@5 0.3092 0.2044 0.2984 0.2803 0.1972 0.3120 0.3388 0.3316
HR@10 0.4194 0.4293 0.4817 0.4204 0.3695 0.4204 0.4610 0.4846
NDCG@5 0.2398 0.2390 0.2184 0.2293 0.1307 0.1729 0.2583 0.2806
NDCG@10 0.2643 0.2433 0.2580 0.2930 0.1713 0.2345 0.2814 0.2939
MRR 0.2294 0.2300 0.2402 0.2203 0.1382 0.2254 0.2581 0.2641
Books HR@1 0.1218 0.1194 0.1177 0.1020 0.0480 0.0853 0.1125 0.1251
HR@5 0.2983 0.3093 0.2764 0.2674 0.1691 0.2465 0.2652 0.3197
HR@10 0.3204 0.3449 0.3781 0.4094 0.3553 0.3582 0.3614 0.4248
NDCG@5 0.1910 0.1980 0.1874 0.1877 0.1031 0.1579 0.1921 0.2097
NDCG@10 0.2573 0.2673 0.2963 0.2673 0.1607 0.2016 0.2415 0.2814
MRR 0.2130 0.2203 0.2599 0.2563 0.1407 0.1832 0.2193 0.2672
Steam HR@1 0.0783 0.0847 0.0960 0.0744 0.0304 0.0641 0.1060 0.1095
HR@5 0.2653 0.2483 0.2665 0.2174 0.1429 0.1966 0.2445 0.2759
HR@10 0.2901 0.2899 0.3170 0.3237 0.2735 0.3218 0.3126 0.3574
NDCG@5 0.1691 0.1335 0.1221 0.1694 0.0915 0.1496 0.1614 0.1735
NDCG@10 0.1739 0.1562 0.1771 0.1884 0.1161 0.1905 0.1809 0.2212
MRR 0.2007 0.1965 0.1882 0.1872 0.1294 0.1528 0.1957 0.2168
Yelp2022 HR@1 0.0771 0.0936 0.1077 0.0724 0.0495 0.0917 0.1005 0.0989
HR@5 0.2766 0.2355 0.2687 0.2211 0.1513 0.2405 0.2395 0.2869
HR@10 0.3108 0.3362 0.3760 0.3228 0.2996 0.3395 0.3317 0.3966
NDCG@5 0.1733 0.1823 0.1995 0.1931 0.1201 0.1397 0.2067 0.2078
NDCG@10 0.2034 0.2164 0.2234 0.2127 0.1225 0.1861 0.2145 0.2277
MRR 0.2093 0.2029 0.1842 0.1980 0.1356 0.1578 0.1906 0.2100
5.2 Result Analysis
From the approaches perspective, IKGR shows new state-of-the-art performances especially on the Search dataset with the largest knowledge gap. On the other hand, across these datasets, IKGR performs better than implicit intent approach (KGIN) or intent unaware LLM approaches (LLMRec), demonstrating the gain from explicit entity node augmentation.
5.3 Ablation Studies
To investigate the influence of each component over the enterprise search data, we conduct a list of experiments that drop each component in the proposed pipeline and compare with the original pipeline on the Search and Beauty dataset. Results are presented in Table 3.
Table 3: Ablation Study
Dataset Versions HR@1 HR@5 HR@10 NDCG@5 NDCG@10 MRR
Search Opt1. Int. Prior GNN 0.0086 0.0202 0.0267 0.0151 0.0172 0.0153
Opt2. Vanilla GNN 0.0078 0.0184 0.048 0.0132 0.0152 0.0130
Opt3. Vanilla Trans. 0.0082 0.0193 0.0256 0.0140 0.0164 0.0139
Opt4. Vanilla Scoring 0.0084 0.0199 0.0258 0.0142 0.0161 0.0137
No Related Intent 0.0077 0.0185 0.0254 0.0147 0.0163 0.0134
No Intent 0.0073 0.0175 0.0240 0.0131 0.0150 0.0125
Beauty Opt1. Int. Prior GNN 0.1369 0.3316 0.4846 0.2806 0.2939 0.2641
Opt2. Vanilla GNN 0.1294 0.3201 0.4299 0.2537 0.2674 0.2502
Opt3. Vanilla Trans. 0.1311 0.3193 0.4272 0.2519 0.2566 0.2439
Opt4. Vanilla Scoring 0.1332 0.3214 0.4249 0.2555 0.2579 0.2522
No Related Intent 0.1266 0.3284 0.4462 0.2643 0.2741 0.2536
No Intent 0.1183 0.2984 0.4093 0.2463 0.2453 0.2399
In Table 3, Opt1. Int. Prior GNN means the full IKGR version without component dropping. Opt2. Vanilla GNN means to use a plain GNN to make predictions over the generated user-intent-item graph. Opt3. Vanilla Trans. means removing both GNN and intent-aware scoring function, and only use a plain graph translation layer. Opt4. Vanilla Scoring means removing the intent prior scoring. These four options correspond to four candidates in modeling the intent graph, all detailed in Appendix B. No Related Intent means dropping the second round of related intent retrieval using RAG, and only exact intent edges are presented in the graph, without the related intent edges. No Intent drops all intent nodes and simply use the GNN of IKGR to predict over user-item graph.
As observed in the results shown in Table 3, the relative contributions of each components of IKGR can be estimated. The most significant observation is that the intent edges augmentation steps as a whole helps to densify the graph almost twice, as reflected by the #IntEdges and #Inter in Table 1, and the MRR performance correspondingly improved 22% (0.0125 to 0.0153 in Search dataset). This confirms that adding intent relations to the graph improves the IKGR performance while densifying the graph with meaningful knowledge connectivities.
5.4 Cold Start Metrics
We use the books dataset to evaluate the cold start setting. The evaluation set is chosen to be a subset of edges whose end nodes both have less than or equal to 3 interactions (node degrees). The results are presented in Table 4. As can be seen from the table, IKGR achieves better performance than other methods on the tail edge set, validating its effectiveness of graph augmentation in dealing with cold start.
Table 4: Performance comparison over tail edges
Setting HR@10 NDCG@10 MRR
IKGR 0.4085 0.2791 0.2630
ConvNCF $0.2699$ $0.1373$ $0.1382$
FPMC $0.2810$ $0.1427$ $0.1326$
KTUP $0.3114$ $0.2044$ $0.1679$
Table 5: Results on different hyperparameter configurations
Configuration HR@10 NDCG@10 MRR
$k=120$, #conv=1=1 0.4262 0.2835 0.2711
$k=100$, #conv=1=1 0.4248 0.2814 0.2672
$k=80$, #conv=1=1 0.3610 0.2619 0.2520
$k=50$, #conv=1=1 0.3023 0.2205 0.1945
$k=100$, #conv=2=2 0.3735 0.2517 0.2482
5.5 Hyperparameter Sensitiveness
To test how model behaves under different hyperparameters, we computed top-k in the kNN retrieval step of RAG, and number of GNN layers in the model architecture. These experiments are conducted on the Books dataset.
The hyperparameter sensitivity results are shown in Table 5. As seen from the table, the model kNN saturates at k=100, which validates our architecture choice of $k$ that both ensures performance and avoid over lengthy token sequences.
5.6 Impact of Knowledge Base
Table 6: Statistics of extracted intents
Metric Value
Average num entity with KB $15.5$
Average num entity without KB $14.5$
Avg Jaccard Similarity Coefficient $0.892$
To quantify the influence of the knowledge base, we use the enterprise search data to compare the extracted intent sets for two scenarios: the knowledge base appended and dropped. We computed the average number of entity extracted, and the Jaccard Similarity Coefficient ( $J(A,B)=\frac{|A\cap B|}{|A\cup B|}$ ) for the entities under these two settings in Table 6.
The result shows that the average number of intents does not differ much, and the Jaccard Similarity Coefficient is close to 1, meaning the knowledge base has limited impact on the extracted entity set. Our intention with the knowledge base is to serve as a lightweight information source in the event of heavy domain gap, rather than a bottleneck key component. Indeed, we have made the LLM generation task as a simple entity extraction task, hence the impact of missing structured knowledge is minimized.
6 Conclusions
Appendix A Full Prompts And Visualizations
For the enterprise search scenario, the full prompt in the first round of intent generation is provided as follows.
### Task
You are given a short paragraph of description of a dataset. Extract a few entities mentioned in the description that the users may be interested in when they consume this dataset. Acronyms can also be included in your answer, if any. Acronyms connected via underscores can be broken apart and partially used, so as to only include the core part into your answer. Please discard entities that have too general and broad meaning, and only pick those specific to this dataset. Please also discard entities that are over-specific, such as date or seemingly highly specified serial numbers/markers.
Please return answer in the format of python list of string, such as: ["answer 1", "answer 2"]. Only return the python list of string. Do NOT return any other explaination words.
### Dataset Description:
{desc}
### Answer:
The full prompt in the second round of intent linkage densification is provided as follows.
You are given a short paragraph of description of a dataset, and a list of entity candidates. Based on the available entity candidates, which of them are likely to attract the users attention, if the user has proven interest in the dataset? Please include related entities, which may not be diretly mentioned in the dataset description. Please make selections only from the available options given below. Please return answer in the format of python list of string.
### Dataset Description:
{desc}
### Available Options:
{RAG_options}
### Important Notes:
1. You can ONLY choose from the Available Options above. If certain entity is not mentioned in the Available Options, you CANNOT return that entity.
2. Please return as many related entities as possible.
3. Please only return the python list of string. Do NOT return any other explaination words.
### Answer:
For the open source dataset (e.g. Steam), the prompt for description summarization using the open world knowledge agent is as follows.
{item_name} is a game on Steam. Collect necessary information for this game, and write one paragraph summary of why a user is potentially interested in it. Use objective tone.
The above prompt converts the item name into a paragraph of LLM summary. Then LLM is used to extract the intent using a similar prompt as above. An example of extracted intents for steam dataset is as follows.
Origin item name: Half-Life 2
LLM summary: Half-Life 2 is a landmark first-person shooter that remains highly relevant today for its exceptional storytelling, physics-based gameplay, and rich world-building. Players could be interested in it because it offers more than just combat; the games silent protagonist, Gordon Freeman, navigates a deeply immersive narrative that unfolds entirely through gameplay, pulling you into a desperate human resistance against an alien empire. The revolutionary Gravity Gun allows for creative problem-solving and combat by manipulating objects, while the games detailed environments and memorable characters-like the resourceful Alyx Vance-create a compelling and lasting experience that set a new standard for the genre. Ultimately, its a masterclass in game design that is both historically significant and still incredibly fun to play.
Extracted Intent entities: ["storytelling", "physics-based gameplay", "world-building", "combat", "narrative", "human resistance", "alien empire", "gravity gun", "problem-solving", "characters", "alyx vance", "game design"]
Appendix B Intent Prior GNN
Below we specify the details of intent prior GNN and how it improves scoring with intent awareness at test time.
B.1 Knowledge Graph Embedding Translation Formulations
@@
Due to the incomplete nature of KGs, KG completion is often leveraged as a self-supervised learning task, which predicts the missing entity $e_{h}$ or $e_{t}$ for a triplet $(e_{h},e_{t},r)$. To this end, TransE [33], a popular knowledge graph embedding model commonly used for KG completion, enforces a translation in the embedding space: $\mathbf{e}_{h}+\mathbf{e}_{r}\approx\mathbf{e}_{t}$. This embedding translation is achieved via the following training objective: $\min\mathop{\Sigma}\limits_{\mathbf{e}_{h},\mathbf{r},\mathbf{e}_{t}}\parallel\mathbf{e}_{h}+\mathbf{r}-\mathbf{e}_{t}\parallel$.
@@
One issue with TransE embedding is that a single relation type may correspond to multiple head entities or tail entities, leading to significant 1-to-N, N-to-1, and N-to-N issues [34]. As an improvement, TransH [34] learns different representations for an entity conditioned on different relations. It assumes that each relation owns a hyperplane, and the translation between head entity and tail entity is valid only if they are projected to the same hyperplane. It defines an energy score function for a triplet as follows:
$f(\mathbf{e}_{h},\mathbf{e}_{t},\mathbf{r})=\parallel\mathbf{e}_{h}^{\bot}+\mathbf{r}-\mathbf{e}_{t}^{\bot}\parallel$ (4)
@@
$\mathbf{e}_{h}^{\bot}=\mathbf{e}_{h}-\mathbf{w}_{r}^{\mathrm{T}}\mathbf{e}_{h}\mathbf{w}_{r}$ (5)
$\mathbf{e}_{t}^{\bot}=\mathbf{e}_{t}-\mathbf{w}_{r}^{\mathrm{T}}\mathbf{e}_{t}\mathbf{w}_{r}$ (6)
@@
Finally, the training of TransH encourages the discrimination between valid triplets and incorrect ones using margin-based ranking loss. Similar to KTUP [29], we employ:
$\mathcal{L}_{KG}=-\sum_{\begin{subarray}{c}(\mathbf{e}_{h},\mathbf{e}^{+}_{t},\mathbf{r}^{+})\in\mathcal{KG}\\ (\mathbf{e}_{h},\mathbf{e}^{-}_{t},\mathbf{r}^{-})\in\mathcal{KG}^{-}\end{subarray}}\log\sigma[f(\mathbf{e}_{h},\mathbf{e}^{-}_{t},\mathbf{r}^{-})-f(\mathbf{e}_{h},\mathbf{e}^{+}_{t},\mathbf{r}^{+})]$ (7)
@@
Figure 3: The intent prior GNN in IKGR. This module is the knowledge graph convolution layer followed by an embedding translation layer, where the intent translation is encoded as structure prior in the embedding space.
B.2 Intent Prior Injected Translation layers